OpenStudy (anonymous):

evaluate 5a+d-bc= a=5, b=7, c=2, d=4

OpenStudy (anonymous):

5(4)+7-(7)(2)

OpenStudy (anonymous):

\(\Large 5\times 5+4-7\times 2=25+4-14=15\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome